The ph of solution produced when an aqueous sol of stronge acid ph 5 is mixed with equal volumes of other solution having ph  3 is -

Grade:11

1 Answers

Arun
6 years ago

Concentration produced by soln. of pH 5=0.00001mole/ L

Concentration produced by soln. of pH 3=0.001mole/ L

Concentration of H+ ion of resultant solution = ( Concentration of H+ ion produced by soln. of ph5+Concentration of H+ ion  produced by soln. of ph3)/2

  H+=0.00001+0.001/2mole/lt.=0.000505mol/lt

pH= -log[H+]

pH= -log(0.000505)

pH=3.3(ph of resultant solution